OpenStudy (anonymous):

Find the distance between (1, 4) and (3, 9)

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (chihiroasleaf):

the distance between two points \( \left( x_1, y_1 \right)\) and \( \left( x_2,y_2\) is \[\large \sqrt{\left( x_2 -x_1\right)^2 + \left(y_2 -y_1\right)^2}\] can you identify your \( x_1, y_1, x_2, y_2\)?

OpenStudy (anonymous):

@louiseeeeee : thank you but I think this is to find slope :) It has another formula :)

OpenStudy (anonymous):

@chihiroasleaf :Realy right ;) thank you ;) @johnechols17 In there we have : x1=1 x2=3 y1=4 y2=9 Got it ?:)

OpenStudy (anonymous):

@E.ali ow sorry. thank you :))))))))

OpenStudy (anonymous):

No prob ;)